How do I solve the following:
The Future Teachers of America is selling a trail mix for fundraising. The mix will contain 3 times the numbers of raisins as sunflowers seeds. Sunflower seeds cost $4.00 per pound, and raisins cost $1.50 per pound. If the group has $34.00 to spend on the raisins and sunflower seeds, how many pounds of each should they buy?

Answer by scott8148(6628)

"The mix will contain 3 times the numbers of raisins as sunflowers seeds" ___ r = 3s

s(4.00) + r(1.50) = 34.00

substituting ___ s(4.00) + 3s(1.50) = 34.00

8.5s = 34 ___ s = 4

substituting ___ r = 3(4) ___ r = 12